OpenStudy (anonymous):

The function T(n) is defined below. Which of the following are equal to T(8)? Check all that apply. T(n) = 4n - 5 A. 37 B. T(7) + 4 C. 27 D. T(5) + T(3) There could be more than one answer. I know C is one of them.

OpenStudy (anonymous):

what is \(T(7)+4\)?

OpenStudy (anonymous):

I'm not sure. I'm still trying to figure that out.

OpenStudy (anonymous):

how did you get \(T(8)\) ?

OpenStudy (anonymous):

N = 8... right? I think.

OpenStudy (anonymous):

replace \(n\) by \(8\) right

OpenStudy (anonymous):

yea then I multiplied \[4 \times 8 - 5 = 27\]

OpenStudy (anonymous):

ok good now try \[4\times 7-5+4\] and see what you get

OpenStudy (anonymous):

That's 27 as well.

OpenStudy (anonymous):

I don't understand how you got that.